# June, 2024

### v1

@@
The Batalin-Vilkovisky formalism is applied to quantise the ${\cal N}=1$ supersymmetric generalisation of the Freedman-Townsend (FT) model proposed by Lindström and Roček in 1983. This super FT theory describes a non-Abelian tensor multiplet and is known to be classically equivalent to a supersymmetric nonlinear sigma model. Using path integral considerations, we demonstrate that this equivalence holds at the quantum level. A modified Faddeev-Popov procedure is employed to quantise models for the ${\cal N}=2$ tensor multiplet in harmonic superspace.
@@ 1 Introduction
The ${\cal N}=1$ supersymmetric extension of the FT model was introduced by Lindström and Roček in 1983 [37]. [^4] Actually, the authors of [37] gave credit to Cremmer and Ferrara, and Gates and Siegel for this model. Further studies and generalisations of this theory have been given in [38, 39, 40]. This is a non-Abelian extension of the ${\cal N}=1$ tensor multiplet proposed by Siegel [41].
@@
This paper is organised as follows. Section 2 is devoted to: (i) carrying out BV quantisation of the ${\cal N}=1$ super FT model (2.1); and (ii) proving its quantum equivalence to the supersymmetric nonlinear $\sigma$-model (2.11). Section 3 is devoted to covariant quantisation of ${\cal N}=2$ tensor multiplet models. For this goal we we make use of the modified Faddeev-Popov procedure described in [23]. Concluding comments are given in section 4. Rudiments of harmonic superspace are collected in appendix A.
@@ 2 Supersymmetric Freedman-Townsend model
$S_{0}[\chi,\bar{\chi},V]=\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{+}\,\chi^{\alpha}W_{\alpha}+\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{-}\,\bar{\chi}_{\dot{\alpha}}\bar{W}^{\dot{\alpha}}+\frac{1}{2}\text{tr}\int{\rm d}^{4|4}z\,V^{2}~{},$ (2.1)
@@
¯$\bar{\nabla}_{\dot{\alpha}}\chi_{\alpha}=0~{},\qquad\bar{\nabla}_{{\dot{\alpha}}}W_{\alpha}=0~{},\qquad\nabla^{\alpha}W_{\alpha}=\bar{\nabla}_{\dot{\alpha}}\bar{W}^{\dot{\alpha}}~{}.$ (2.2)
Operators $\nabla_{A}=(\nabla_{a},\nabla_{\alpha},\bar{\nabla}^{\dot{\alpha}})$ above are gauge-covariant derivatives and obey the algebra:
$\{\nabla_{\alpha},\nabla_{\beta}\}=0~{},\quad$$\{\nabla_{\alpha},\bar{\nabla}_{\dot{\alpha}}\}=-2{\rm i}\nabla_{\alpha{\dot{\alpha}}}~{},\quad\{\bar{\nabla}_{\dot{\alpha}},\bar{\nabla}_{\dot{\beta}}\}=0~{},$ (2.3a)
$[\nabla_{\alpha},\nabla_{\beta{\dot{\beta}}}]$ $={\rm i}\varepsilon_{\alpha\beta}\bar{W}_{\dot{\alpha}}~{},\qquad[\bar{\nabla}_{\dot{\alpha}},\nabla_{\beta{\dot{\beta}}}]={\rm i}\varepsilon_{{\dot{\alpha}}{\dot{\beta}}}W_{\alpha}~{},$ (2.3b)
$[\nabla_{\alpha{\dot{\alpha}}},\nabla_{\beta{\dot{\beta}}}]$ $=-\frac{1}{2}\varepsilon_{\alpha\beta}\bar{\nabla}_{({\dot{\alpha}}}\bar{W}_{{\dot{\beta}})}-\frac{1}{2}\varepsilon_{{\dot{\alpha}}{\dot{\beta}}}\nabla_{(\alpha}W_{\beta)}~{}.$ (2.3c)
These constraints may be solved [45] in terms of the complex unconstrained prepotential $\Omega=\Omega^{I}T^{I}$ and its Hermitian conjugate $\bar{\Omega}\equiv\Omega^{\dagger}$,
$\nabla_{\alpha}={\rm e}^{-\Omega}D_{\alpha}{\rm e}^{\Omega}~{},\qquad\bar{\nabla}_{\dot{\alpha}}={\rm e}^{\bar{\Omega}}\bar{D}_{\dot{\alpha}}{\rm e}^{-\bar{\Omega}}~{},\qquad{\rm e}^{V}={\rm e}^{\Omega}{\rm e}^{\bar{\Omega}}~{},$ (2.4)
where $D_{A}=(\partial_{a},D_{\alpha},\bar{D}^{\dot{\alpha}})$ are the covariant derivatives of Minkowski superspace. [^7] Our normalisation of the gauge prepotentials in (2.4) differs from [44] and agrees with [46]. The superfields $\bar{\chi}_{\dot{\alpha}}$ and $\bar{W}_{\dot{\alpha}}$ in (2.1) are Hermitian conjugates of $\chi_{\alpha}$ and $W_{\alpha}$, respectively, in the vector representation. The conjugation rules are somewhat different in the chiral representation [46].
@@
$\delta_{L}\chi_{\alpha}=-\frac{{\rm i}}{4}\bar{\nabla}^{2}\nabla_{\alpha}L~{},\qquad L=\bar{L}~{},$ (2.5)
@@ 2.1 Classical equivalence
$S=\int{\rm d}^{4|4}z\,{\mathfrak{F}}(G)~{},\qquad G=\frac{1}{2}(D^{\alpha}\eta_{\alpha}+\bar{D}_{\dot{\alpha}}\bar{\eta}^{\dot{\alpha}})~{},\qquad\bar{D}_{\dot{\beta}}\eta_{\alpha}=0~{}.$ (2.7)
@@
$S=\int{\rm d}^{4|4}z\,{\mathfrak{U}}(\varphi+\bar{\varphi})~{},\qquad\bar{D}_{\dot{\beta}}\varphi=0~{},$ (2.8)
@@
$\frac{\delta_{l}S_{0}[\chi,\bar{\chi},V]}{\delta\chi^{\alpha}}=\frac{1}{2}W_{\alpha}=0\quad\implies\quad V=\log({\rm e}^{\bar{\Phi}}{\rm e}^{\Phi}\big{)}~{},\qquad\bar{D}_{\dot{\alpha}}\Phi=0~{}.$ (2.10)
@@
One of the goals of this section, which we pursue below, is to prove that this equivalence extends to the quantum level. Specifically, our goal is to demonstrate that the partition functions for models (2.1) and (2.11) coincide.
@@ 2.2 Quantisation via Batalin-Vilkovisky formalism
$\delta_{L}\chi^{\alpha I}(z)$ $=\int{\rm d}^{4|4}z\,{R}^{\alpha I\,J}(z,z^{\prime})L^{J}(z^{\prime})\quad\implies\quad{R}^{\alpha I\,J}(z,z^{\prime})=-\frac{{\rm i}}{4}\bar{\nabla}^{2}\nabla^{\alpha}\delta^{8}(z-z^{\prime})\delta^{IJ}~{},$ (2.14a)
$\delta_{L}\bar{\chi}_{\dot{\alpha}}^{I}(z)$ $=\int{\rm d}^{4|4}z\,{R}_{\dot{\alpha}}^{I\,J}(z,z^{\prime})L^{J}(z^{\prime})\quad\implies\quad{R}_{\dot{\alpha}}^{I\,J}(z,z^{\prime})=\frac{{\rm i}}{4}\nabla^{2}\bar{\nabla}_{\dot{\alpha}}\delta^{8}(z-z^{\prime})\delta^{IJ}~{}.$ (2.14b)
@@
$\delta_{\sigma}L^{I}(z)$ $=\int{\rm d}^{4|2}z_{+}\,Z^{I\,J}(z,z^{\prime})\sigma(z^{\prime})+\int{\rm d}^{4|2}z_{-}\,\bar{Z}^{I\,J}(z,z^{\prime})\bar{\sigma}(z^{\prime})~{},$ (2.16a)
@@
$Z=\int\big{[}\mathscr{D}\mbox{\boldmath$\Phi$}^{\underline{A}}\big{]}\big{[}\mathscr{D}\mbox{\boldmath$\Phi$}^{*}_{\underline{A}}\big{]}\,\mu\,\delta\Big{[}\mbox{\boldmath$\Phi$}^{*}_{\underline{A}}-\frac{\delta\Psi}{\delta{\mbox{\boldmath$\Phi$}}^{\underline{A}}}\Big{]}\exp\Big{(}\frac{{\rm i}}{\hbar}S[\mbox{\boldmath$\Phi$},\mbox{\boldmath$\Phi$}^{*}]\Big{)}~{},$ (2.19)
@@
$S[\mbox{\boldmath$\Phi$}_{\text{min}},{\mbox{\boldmath$\Phi$}}^{*}_{\text{min}}]$ $=S_{0}[\chi,\bar{\chi},V]+\text{tr}\int{\rm d}^{4|4}z\,\Big{\{}{\rm i}\big{(}\nabla^{\alpha}\chi_{\alpha}^{*}-\bar{\nabla}_{\dot{\alpha}}\bar{\chi}^{*{\dot{\alpha}}}\big{)}C+C^{*}(\eta+\bar{\eta})\Big{\}}$
$\phantom{=}+{\rm i}\,\text{tr}\int{\rm d}^{4|2}z_{+}\,\big{[}\chi^{*\alpha},\chi_{\alpha}^{*}\big{]}\eta-{\rm i}\,\text{tr}\int{\rm d}^{4|2}z_{-}\,\big{[}\bar{\chi}^{*}_{\dot{\alpha}},\bar{\chi}^{{\dot{\alpha}}*}\big{]}\bar{\eta}~{},$ (2.22)
@@
$S[\mbox{\boldmath$\Phi$},\mbox{\boldmath$\Phi$}^{*}]$ $=S[\mbox{\boldmath$\Phi$}_{\text{min}},{\mbox{\boldmath$\Phi$}}^{*}_{\text{min}}]+\text{tr}\int{\rm d}^{4|4}z\,\Big{\{}\tilde{C}^{*}\pi+\mbox{\boldmath$\eta$}^{*}\bar{\mbox{\boldmath$\pi$}}+\bar{\mbox{\boldmath$\eta$}}^{*}\mbox{\boldmath$\pi$}+\eta^{\prime*}\bar{\pi}^{\prime}+\bar{\eta}^{\prime*}\pi^{\prime}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\Lambda^{*}\bar{\Pi}+\bar{\Lambda}^{*}\Pi+\Lambda^{\prime*}\bar{\Pi}^{\prime}+\bar{\Lambda}^{*}\Pi^{\prime}\Big{\}}$ (2.24)
@@
$\Psi(\mbox{\boldmath$\Phi$})$ $=\text{tr}\int{\rm d}^{4|4}z\,\Big{\{}\frac{{\rm i}}{2}\tilde{C}\big{(}\nabla^{\alpha}\chi_{\alpha}-\bar{\nabla}_{\dot{\alpha}}\bar{\chi}^{\dot{\alpha}}\big{)}+C(\mbox{\boldmath$\eta$}+\bar{\mbox{\boldmath$\eta$}})+\tilde{C}(\eta^{\prime}+\bar{\eta}^{\prime})-\frac{1}{2m}\tilde{C}\pi\Big{\}}$
$\phantom{=}\quad+\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{+}\,\Big{\{}-\frac{1}{n}\big{(}{\mbox{\boldmath$\eta$}}\pi^{\prime}+\mbox{\boldmath$\pi$}\eta^{\prime}\big{)}+\Lambda\Pi+\Lambda^{\prime}\Pi^{\prime}\Big{\}}$
$\phantom{=}\quad+\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{-}\,\Big{\{}-\frac{1}{n}\big{(}\bar{\mbox{\boldmath$\eta$}}\bar{\pi}^{\prime}+\bar{\mbox{\boldmath$\pi$}}\bar{\eta}^{\prime}\big{)}+\bar{\Lambda}\bar{\Pi}+\bar{\Lambda}^{\prime}\bar{\Pi}^{\prime}\Big{\}}~{},$ (2.26)
@@
$S[\mbox{\boldmath$\Phi$},{\mbox{\boldmath$\Phi$}}^{*}]$ $=S_{0}[\varphi]+\text{tr}\int{\rm d}^{4|4}z\,\Big{\{}\frac{1}{8}\tilde{C}\big{(}\nabla^{\alpha}\bar{\nabla}^{2}\nabla_{\alpha}+\bar{\nabla}_{\dot{\alpha}}\nabla^{2}\bar{\nabla}^{\dot{\alpha}}\big{)}C+\bar{\eta}\mbox{\boldmath$\eta$}+\bar{\mbox{\boldmath$\eta$}}\eta-\bar{\Pi}\Pi-\bar{\Pi}^{\prime}\Pi^{\prime}$
$\qquad\qquad\qquad+\Big{(}\frac{{\rm i}}{2}\big{(}\nabla^{\alpha}\chi_{\alpha}-\bar{\nabla}_{\dot{\alpha}}\bar{\chi}^{\dot{\alpha}}\big{)}+(\eta^{\prime}+\bar{\eta}^{\prime})-\frac{1}{2m}\pi\Big{)}\pi$
$\qquad\qquad\qquad+\Big{(}-\frac{1}{2n}\pi^{\prime}-\frac{1}{4}\bar{\nabla}^{2}C\Big{)}\bar{\mbox{\boldmath$\pi$}}+\Big{(}-\frac{1}{2n}\bar{\pi}^{\prime}-\frac{1}{4}\nabla^{2}C\Big{)}\mbox{\boldmath$\pi$}$
$\qquad\qquad\qquad+\Big{(}\frac{1}{2n}\mbox{\boldmath$\pi$}-\frac{1}{4}\bar{\nabla}^{2}\tilde{C}\Big{)}\bar{\pi}^{\prime}+\Big{(}\frac{1}{2n}\bar{\mbox{\boldmath$\pi$}}-\frac{1}{4}\nabla^{2}\tilde{C}\Big{)}\pi^{\prime}\Big{\}}$
$\phantom{=}\quad-\frac{{\rm i}}{64}\,\text{tr}\int{\rm d}^{4|2}z_{+}\,\big{[}\bar{\nabla}^{2}\nabla^{\alpha}\tilde{C},\bar{\nabla}^{2}\nabla_{\alpha}\tilde{C}\big{]}\eta+\frac{{\rm i}}{64}\text{tr}\int{\rm d}^{4|2}z_{-}\,\big{[}\nabla^{2}\bar{\nabla}_{\dot{\alpha}}\tilde{C},\nabla^{2}\bar{\nabla}^{\dot{\alpha}}\tilde{C}\big{]}\bar{\eta}~{}.$ (2.27)

### v2

@@
The Batalin-Vilkovisky formalism is applied to quantise the ${\cal N}=1$ supersymmetric generalisation of the Freedman-Townsend (FT) model, which was proposed by Lindström and Roček in 1983 in Minkowski superspace and is lifted to a supergravity background in this paper. This super FT theory describes a non-Abelian tensor multiplet and is known to be classically equivalent to a supersymmetric nonlinear sigma model. Using path integral considerations, we demonstrate that this equivalence holds at the quantum level in the sense that the quantum supercurrents in the two theories coincide. A modified Faddeev-Popov procedure is employed to quantise models for the ${\cal N}=2$ tensor multiplet in harmonic superspace. The obtained results agree with those derived by applying the Batalin-Vilkovisky scheme within the harmonic superspace setting.
@@ 1 Introduction
The ${\cal N}=1$ supersymmetric extension of the FT model was introduced by Lindström and Roček in 1983 [37]. [^4] Actually, the authors of [37] gave credit to Cremmer and Ferrara, and Gates and Siegel for this model. Further studies and generalisations of this theory have been given in [38, 39, 40]. This is a non-Abelian extension of the ${\cal N}=1$ tensor multiplet proposed by Siegel [41]. In this paper we will couple the ${\cal N}=1$ super FT model to supergravity.
@@
This paper is organised as follows. Section 2 is devoted to: (i) coupling the super FT model to old minimal supergravity; (ii) carrying out BV quantisation of the ${\cal N}=1$ super FT model (2.1); and (iii) proving its quantum equivalence to the locally supersymmetric nonlinear $\sigma$-model (2.13). Section 3 is devoted to covariant quantisation of ${\cal N}=2$ tensor multiplet models. For this goal we we make use of the modified Faddeev-Popov procedure described in [23]. Concluding comments are given in section 4. Rudiments of harmonic superspace are collected in appendix A. BV quantisation of the ${\cal N}=2$ Abelian vector and tensor multiplets is carried out in appendix B.
@@ 2 Supersymmetric Freedman-Townsend model
$S_{0}[\chi,\bar{\chi},V]=\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{+}\,\mathcal{E}\,\chi^{\alpha}W_{\alpha}+\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{-}\,\bar{\mathcal{E}}\,\bar{\chi}_{\dot{\alpha}}\bar{W}^{\dot{\alpha}}+\frac{1}{2}\text{tr}\int{\rm d}^{4|4}z\,E\,V^{2}~{},$ (2.1)
@@
¯$\bar{\nabla}_{\dot{\alpha}}\chi_{\alpha}=0~{},\qquad\bar{\nabla}_{{\dot{\alpha}}}W_{\alpha}=0~{},\qquad\nabla^{\alpha}W_{\alpha}=\bar{\nabla}_{\dot{\alpha}}\bar{W}^{\dot{\alpha}}~{},$ (2.2)
and $E^{-1}={\rm Ber}(E_{A}{}^{M})$ is the full superspace measure, while $\mathcal{E}$ denotes the chiral density. Operators $\nabla_{A}=(\nabla_{a},\nabla_{\alpha},\bar{\nabla}^{\dot{\alpha}})$ above are gauge-covariant derivatives, $\nabla_{A}=E_{A}{}^{M}\partial_{M}+\frac{1}{2}\Omega_{A}{}^{bc}M_{bc}+{\rm i}\,\Gamma_{A}{}^{I}T^{I}$, with $M_{bc}$ the Lorentz generators. The covariant derivatives obey the algebra:
$\{\nabla_{\alpha},\nabla_{\beta}\}$ $=-4\bar{R}M_{\alpha\beta}~{},\quad\{\nabla_{\alpha},\bar{\nabla}_{\dot{\alpha}}\}=-2{\rm i}\nabla_{\alpha{\dot{\alpha}}}~{},\quad\{\bar{\nabla}_{\dot{\alpha}},\bar{\nabla}_{\dot{\beta}}\}=4R\bar{M}_{{\dot{\alpha}}{\dot{\beta}}}~{},$ (2.3a)
[¯$[\bar{\nabla}_{\dot{\alpha}},\nabla_{\beta{\dot{\beta}}}]$ $=-{\rm i}\varepsilon_{{\dot{\alpha}}{\dot{\beta}}}\Big{(}R\nabla_{\beta}+G_{\beta}{}^{{\dot{\gamma}}}\bar{\nabla}_{\dot{\gamma}}-\bar{\nabla}^{\dot{\gamma}}G_{\beta}{}^{{\dot{\delta}}}\bar{M}_{{\dot{\gamma}}{\dot{\delta}}}+2W_{\beta}{}^{\gamma\delta}M_{\gamma\delta}-{W}_{\beta}\Big{)}-{\rm i}\nabla_{\beta}R\bar{M}_{{\dot{\alpha}}{\dot{\beta}}}~{},$ (2.3b)
$[{\nabla}_{\alpha},\nabla_{\beta{\dot{\beta}}}]$ $={\rm i}\varepsilon_{\alpha\beta}\Big{(}\bar{R}\bar{\nabla}_{\dot{\beta}}+G^{\gamma}{}_{{\dot{\beta}}}{\nabla}_{\gamma}-{\nabla}^{\gamma}G^{\delta}{}_{\dot{\beta}}\bar{M}_{{\dot{\gamma}}{\dot{\delta}}}\bar{M}_{{\dot{\gamma}}{\dot{\delta}}}+2\bar{W}_{\dot{\beta}}{}^{{\dot{\gamma}}{\dot{\delta}}}\bar{M}_{{\dot{\gamma}}{\dot{\delta}}}+\bar{W}_{\dot{\beta}}\Big{)}+{\rm i}\bar{\nabla}_{{\dot{\beta}}}\bar{R}\bar{M}_{{\dot{\alpha}}{\dot{\beta}}}~{}.$ (2.3c)
Here the torsion tensors $R$, $G_{a}={\bar{G}}_{a}$ and $W_{\alpha\beta\gamma}=W_{(\alpha\beta\gamma)}$ satisfy the Bianchi identities:
¯$\bar{\nabla}_{\dot{\alpha}}R=0~{},~{}~{}~{}~{}~{}~{}\bar{\nabla}_{\dot{\alpha}}W_{\alpha\beta\gamma}=0~{},$ (2.4a)
¯$\bar{\nabla}^{\dot{\gamma}}G_{\alpha{\dot{\gamma}}}=\nabla_{\alpha}R~{},~{}~{}~{}~{}~{}~{}\nabla^{\gamma}W_{\alpha\beta\gamma}={\rm i}\,\nabla_{(\alpha}{}^{\dot{\gamma}}G_{\beta){\dot{\gamma}}}~{}.$ (2.4b)
In this supergeometry integrals over the full superspace and its chiral subspace are related via:
$\int{\rm d}^{4|4}z\,E\,U=-\frac{1}{4}\int{\rm d}^{4|2}z_{+}\,\,{\cal E}\,(\bar{\nabla}^{2}-4R)U~{}.$ (2.5)
Constraints (2.3) may be solved [45] in terms of the complex unconstrained prepotential $\Omega=\Omega^{I}T^{I}$ and its Hermitian conjugate $\bar{\Omega}\equiv\Omega^{\dagger}$,
$\nabla_{\alpha}={\rm e}^{-\Omega}\mathfrak{D}_{\alpha}{\rm e}^{\Omega}~{},\qquad\bar{\nabla}_{\dot{\alpha}}={\rm e}^{\bar{\Omega}}\bar{\mathfrak{D}}_{\dot{\alpha}}{\rm e}^{-\bar{\Omega}}~{},\qquad{\rm e}^{V}={\rm e}^{\Omega}{\rm e}^{\bar{\Omega}}~{},$ (2.6)
where $\mathfrak{D}_{A}=(\mathfrak{D}_{a},\mathfrak{D}_{\alpha},\bar{\mathfrak{D}}^{\dot{\alpha}})$ are the covariant derivatives of the GWZ superspace [46], see e.g. [44, 47] for a review. [^7] Relation of the GWZ superspace geometry to other ${\cal N}=1$ supergravity formulations in spelled out in [48]. Our normalisation of the gauge prepotentials in (2.6) differs from [44] and agrees with [47]. The superfields $\bar{\chi}_{\dot{\alpha}}$ and $\bar{W}_{\dot{\alpha}}$ in (2.1) are Hermitian conjugates of $\chi_{\alpha}$ and $W_{\alpha}$, respectively, in the vector representation. The conjugation rules are somewhat different in the chiral representation [47].
@@
$\delta_{L}\chi_{\alpha}=-\frac{{\rm i}}{4}(\bar{\nabla}^{2}-4R)\nabla_{\alpha}L~{},\qquad L=\bar{L}~{},$ (2.7)
@@ 2.1 Classical equivalence
$S=\int{\rm d}^{4|4}z\,E\,{\mathfrak{F}}(G)~{},\qquad G=\frac{1}{2}(\mathfrak{D}^{\alpha}\eta_{\alpha}+\bar{\mathfrak{D}}_{\dot{\alpha}}\bar{\eta}^{\dot{\alpha}})~{},\qquad\bar{\mathfrak{D}}_{\dot{\beta}}\eta_{\alpha}=0~{}.$ (2.9)
@@
$S=\int{\rm d}^{4|4}z\,E\,{\mathfrak{U}}(\varphi+\bar{\varphi})~{},\qquad\bar{\mathfrak{D}}_{\dot{\beta}}\varphi=0~{},$ (2.10)
@@
$\frac{\delta_{l}S_{0}[\chi,\bar{\chi},V]}{\delta\chi^{\alpha}}=\frac{1}{2}W_{\alpha}=0\quad\implies\quad V=\log({\rm e}^{\bar{\Phi}}{\rm e}^{\Phi}\big{)}~{},\qquad\bar{\mathfrak{D}}_{\dot{\alpha}}\Phi=0~{}.$ (2.12)
@@
One of the goals of this section, which we pursue below, is to prove that this equivalence extends to the quantum level. Specifically, our goal is to demonstrate that the partition functions for models (2.1) and (2.13) differ by a topological invariant. This means that the quantum supercurrents in these theories coincide.
@@ 2.2 Quantisation via Batalin-Vilkovisky formalism
$\delta_{L}\chi^{\alpha I}(z)$ $=\int{\rm d}^{4|4}z\,E\,{R}^{\alpha I\,J}(z,z^{\prime})L^{J}(z^{\prime})\quad\implies\quad{R}^{\alpha I\,J}(z,z^{\prime})=-\frac{{\rm i}}{4}(\bar{\nabla}^{2}-4R)\nabla^{\alpha}\delta^{8}(z,z^{\prime})\delta^{IJ}~{},$ (2.16a)
$\delta_{L}\bar{\chi}_{\dot{\alpha}}^{I}(z)$ $=\int{\rm d}^{4|4}z\,E\,{R}_{\dot{\alpha}}^{I\,J}(z,z^{\prime})L^{J}(z^{\prime})\quad\implies\quad{R}_{\dot{\alpha}}^{I\,J}(z,z^{\prime})=\frac{{\rm i}}{4}(\nabla^{2}-4\bar{R})\bar{\nabla}_{\dot{\alpha}}\delta^{8}(z,z^{\prime})\delta^{IJ}~{}.$ (2.16b)
@@
$\delta_{\sigma}L^{I}(z)$ $=\int{\rm d}^{4|2}z_{+}\,{\cal E}\,Z^{I\,J}(z,z^{\prime})\sigma^{J}(z^{\prime})+\int{\rm d}^{4|2}z_{-}\,\bar{{\cal E}}\,\bar{Z}^{I\,J}(z,z^{\prime})\bar{\sigma}^{J}(z^{\prime})~{},$ (2.18a)
@@
$Z=\int\big{[}\mathscr{D}\mbox{\boldmath$\Phi$}^{\underline{A}}\big{]}\big{[}\mathscr{D}\mbox{\boldmath$\Phi$}^{*}_{\underline{A}}\big{]}\,\mu\,\delta\Big{[}\mbox{\boldmath$\Phi$}^{*}_{\underline{A}}-\frac{\delta\Psi}{\delta{\mbox{\boldmath$\Phi$}}^{\underline{A}}}\Big{]}\exp\Big{(}{\rm i}S[\mbox{\boldmath$\Phi$},\mbox{\boldmath$\Phi$}^{*}]\Big{)}~{},$ (2.21)
@@
$S[\mbox{\boldmath$\Phi$}_{\text{min}},{\mbox{\boldmath$\Phi$}}^{*}_{\text{min}}]$ $=S_{0}[\chi,\bar{\chi},V]+\text{tr}\int{\rm d}^{4|4}z\,E\,\Big{\{}{\rm i}\big{(}\nabla^{\alpha}\chi_{\alpha}^{*}-\bar{\nabla}_{\dot{\alpha}}\bar{\chi}^{*{\dot{\alpha}}}\big{)}C+C^{*}(\eta+\bar{\eta})\Big{\}}$
$\phantom{=}+{\rm i}\,\text{tr}\int{\rm d}^{4|2}z_{+}\,{\cal E}\,\big{[}\chi^{*\alpha},\chi_{\alpha}^{*}\big{]}\eta-{\rm i}\,\text{tr}\int{\rm d}^{4|2}z_{-}\,\bar{{\cal E}}\,\big{[}\bar{\chi}^{*}_{\dot{\alpha}},\bar{\chi}^{{\dot{\alpha}}*}\big{]}\bar{\eta}~{},$ (2.24)
@@
$S[\mbox{\boldmath$\Phi$},\mbox{\boldmath$\Phi$}^{*}]$ $=S[\mbox{\boldmath$\Phi$}_{\text{min}},{\mbox{\boldmath$\Phi$}}^{*}_{\text{min}}]+\text{tr}\int{\rm d}^{4|4}z\,E\,\Big{\{}\tilde{C}^{*}\pi+\frac{1}{2}\Big{[}(\mbox{\boldmath$\eta$}^{*}+\bar{\mbox{\boldmath$\pi$}})^{2}+(\bar{\mbox{\boldmath$\eta$}}^{*}+\mbox{\boldmath$\pi$})^{2}+(\eta^{\prime*}+\bar{\pi}^{\prime})^{2}$
$\qquad+(\bar{\eta}^{\prime*}+\pi^{\prime})^{2}+(\Lambda^{*}+\bar{\Pi})^{2}+(\bar{\Lambda}^{*}+\Pi)^{2}+(\Lambda^{\prime*}+\bar{\Pi}^{\prime})^{2}+(\bar{\Lambda}^{*}+\Pi^{\prime})^{2}\Big{]}\Big{\}}$ (2.26)
@@
$\Psi(\mbox{\boldmath$\Phi$})$ $=\text{tr}\int{\rm d}^{4|4}z\,E\,\Big{\{}\frac{{\rm i}}{2}\tilde{C}\big{(}\nabla^{\alpha}\chi_{\alpha}-\bar{\nabla}_{\dot{\alpha}}\bar{\chi}^{\dot{\alpha}}\big{)}+C(\mbox{\boldmath$\eta$}+\bar{\mbox{\boldmath$\eta$}})+\tilde{C}(\eta^{\prime}+\bar{\eta}^{\prime})-\frac{1}{2m}\tilde{C}\pi\Big{\}}$
$\phantom{=}\quad+\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{+}\,{\cal E}\,\Big{\{}-\frac{1}{n}\big{(}{\mbox{\boldmath$\eta$}}\pi^{\prime}+\mbox{\boldmath$\pi$}\eta^{\prime}\big{)}+\Lambda\Pi+\Lambda^{\prime}\Pi^{\prime}\Big{\}}$
$\phantom{=}\quad+\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{-}\,\bar{{\cal E}}\,\Big{\{}-\frac{1}{n}\big{(}\bar{\mbox{\boldmath$\eta$}}\bar{\pi}^{\prime}+\bar{\mbox{\boldmath$\pi$}}\bar{\eta}^{\prime}\big{)}+\bar{\Lambda}\bar{\Pi}+\bar{\Lambda}^{\prime}\bar{\Pi}^{\prime}\Big{\}}~{},$ (2.28)
@@
$S[\mbox{\boldmath$\Phi$},{\mbox{\boldmath$\Phi$}}^{*}]$ $=S_{0}[\varphi]+\text{tr}\int{\rm d}^{4|4}z\,E\,\Big{\{}\frac{1}{8}\tilde{C}\big{(}\nabla^{\alpha}(\bar{\nabla}^{2}-4R)\nabla_{\alpha}+\bar{\nabla}_{\dot{\alpha}}(\nabla^{2}-4\bar{R})\bar{\nabla}^{\dot{\alpha}}\big{)}C+\frac{1}{2}(\bar{\eta}+\mbox{\boldmath$\eta$})^{2}$
$\phantom{=}\quad+\frac{1}{2}(\bar{\mbox{\boldmath$\eta$}}+\eta)^{2}-\frac{1}{2}(\bar{\Pi}+\Pi)^{2}-\frac{1}{2}(\bar{\Pi}^{\prime}+\Pi^{\prime})^{2}+\Big{(}\frac{{\rm i}}{2}\big{(}\nabla^{\alpha}\chi_{\alpha}-\bar{\nabla}_{\dot{\alpha}}\bar{\chi}^{\dot{\alpha}}\big{)}+(\eta^{\prime}+\bar{\eta}^{\prime})-\frac{1}{2m}\pi\Big{)}\pi$
$\phantom{=}\quad+\frac{1}{2}\Big{(}-\frac{1}{2n}\pi^{\prime}-\frac{1}{4}(\bar{\nabla}^{2}-4R)C+\bar{\mbox{\boldmath$\pi$}}\Big{)}^{2}+\frac{1}{2}\Big{(}-\frac{1}{2n}\bar{\pi}^{\prime}-\frac{1}{4}(\nabla^{2}-4\bar{R})C+\mbox{\boldmath$\pi$}\Big{)}^{2}$
$\phantom{=}\quad+\frac{1}{2}\Big{(}\frac{1}{2n}\mbox{\boldmath$\pi$}-\frac{1}{4}(\bar{\nabla}^{2}-4R)\tilde{C}+\bar{\pi}^{\prime}\Big{)}^{2}+\frac{1}{2}\Big{(}\frac{1}{2n}\bar{\mbox{\boldmath$\pi$}}-\frac{1}{4}(\nabla^{2}-4\bar{R})\tilde{C}+\pi^{\prime}\Big{)}^{2}\Big{\}}$
$\phantom{=}\quad-\frac{{\rm i}}{64}\,\text{tr}\int{\rm d}^{4|2}z_{+}\,{\cal E}\,\big{[}(\bar{\nabla}^{2}-4R)\nabla^{\alpha}\tilde{C},(\bar{\nabla}^{2}-4R)\nabla_{\alpha}\tilde{C}\big{]}\eta$
$\phantom{=}\quad+\frac{{\rm i}}{64}\text{tr}\int{\rm d}^{4|2}z_{-}\,\bar{{\cal E}}\,\big{[}(\nabla^{2}-4\bar{R})\bar{\nabla}_{\dot{\alpha}}\tilde{C},(\nabla^{2}-4\bar{R})\bar{\nabla}^{\dot{\alpha}}\tilde{C}\big{]}\bar{\eta}~{}.$ (2.29)